solve for x & y when x/3 + y/4= 19/12 AND x/2 - y/5= -63/80
Answer by amarjeeth123(574)

x/3 + y/4= 19/12
Multiplying on both sides by 12 we get,
4x+3y=19.........equation 1
x/2 - y/5= -63/80
Multiplying on both sides by 80 we get,
40x-16y=-63......equation 2
Multiplying equation 1 by 10 we get,
40x+30y=190......equation 3
Subtracting equation 2 from equation 3 we get,
40x+30y-40x+16y=190+63
46y=253
y=253/46.
x=(19-759/46)/4=115/184.